# Exploring the parameter dependence of atomic minima with implicit differentiation

### v1

@@ II.5 Backpropagation of parameter variations for inverse design problems
Inverse design aims to produce materials with specified (desirable) properties through inverting structure-property relationships. However, this typically requires high-throughput searches which necessarily cannot afford to perform atomistic simulations of e.g. defect structures and mechanisms to predict mechanical properties. The implicit derivative can be used to make a first step towards gradient-led inverse design, allowing us to find interatomic parameters ${{\bm{\Theta}}}$ that stabilise some atomic configuration ${\bf X}^{*}$ observed in an ab initio-accurate simulation e.g. DFT or hybrid DFT-ML [4]. With ${\bf X}^{*}_{{{\bm{\Theta}}}}$ being the local minimum found when minimizing $U({\bf X},;{{\bm{\Theta}}})$ starting from ${\bf X}^{*}$, we can write an implicit loss function
$L({{\bm{\Theta}}})=\frac{1}{2}\,\big{[}{\bf X}^{*}_{{{\bm{\Theta}}}}-{\bf X}^{*}\big{]}:\big{[}{\bf X}^{*}_{{{\bm{\Theta}}}}-{\bf X}^{*}\big{]}$ (9)
where we use the notation :::: to indicates summation over the $N$ atomic sites and $3$ spatial indices. The implicit derivative is necessary to compute the derivative of the loss via the chain rule:
$\nabla_{{{\bm{\Theta}}}}L({{\bm{\Theta}}})=\nabla_{{{\bm{\Theta}}}}{\bf X}^{*}_{{{\bm{\Theta}}}}:\big{[}{\bf X}^{*}_{{{\bm{\Theta}}}}-{\bf X}^{*}\big{]}\in\mathbb{R}^{N_{D}}.$ (10)
An immediate application of (9) is the ability to ‘fine-tune’, or retrain[4], interatomic potentials to reproduce important DFT minima. Fine-tuning has gained increasing interest following the rise of ‘foundation model’ machine learning potentials [13], which are beyond the scope of this study. In practice, one typically includes the loss against the original training database to regularize the fine-tuning fit. However, we have found that minimizing (9) in practice produces only very small perturbations to the final potential.
The ability to fine-tune interactions is of particular relevance for low energy structures such as dislocation lines[8, 4], which typically require careful weighting in potential fitting[13]. In section II.8, we use implicit loss minimization to find solute substitutions which induce ‘hard’ screw dislocation core reconstruction in tungsten, and fine-tune a tungsten-beryllium potential to correctly reproduce ab initio observations [4].
@@ II.6 Prediction of the total relaxed energy in classical potentials with automatic differentiation
The binary Lennard-Jones potential is a classical model for nanoclusters and glassy systems [5]. The model is defined by six parameters ${{\bm{\Theta}}}=[\epsilon^{\rm LJ}_{\rm AA},\epsilon^{\rm LJ}_{\rm AB},\epsilon^{\rm LJ}_{\rm BB},\sigma_{\rm AA},\sigma_{\rm AB},\sigma_{\rm BB}]$, with a total energy
@@
To simplify the problem, we set $\epsilon^{\rm LJ}_{\rm AA}=\epsilon^{\rm LJ}_{\rm AB}=\epsilon^{\rm LJ}_{\rm BB}$ and $\sigma_{\rm AA}=\sigma_{\rm BB}=1$, leaving $\Theta=\sigma_{\rm AB}$ as the only varying parameter in this example. When $\sigma_{\mathrm{AB}}=1$, all atoms are identical and the system is a unary fcc lattice; we additionally remove one atom to form a vacancy and promote additional deformation. When $\sigma_{\mathrm{AB}}\neq 1$ the system becomes a random fcc binary alloy, with lattice distortion in the bulk and around the vacancy (see Fig. 2a and 2b). Figure 2c shows the inhomogeneous implicit derivative around $\sigma_{AB}=1$ that gives an excellent prediction of the total energy and lattice distortion for $\sigma_{AB}\in[0.95,1.05]$, with mild disagreement as $|\sigma_{AB}-1|$ grows. At constant volume both constant and homogeneous approximations are equivalent and predict no structural change, with significantly higher errors. However, in the remainder of this paper, we focus on modern machine learning potentials.
@@ II.7 Prediction of defect formation energies and volumes with machine learning potentials
Almost all modern interatomic models use high-dimensional regression techniques from the machine learning community[29, 30, 8, 13]. A common first step is to represent atomic environments as $N_{D}$ per-atom descriptor functions $D_{l}(\tilde{\bf X},{\bf C},i)$, $l\in[1,N_{D}]$ which represent the atomic environment around an atom of index $i$. In practice, the descriptor functions also have species-dependent hyperparameters which must also be tuned, but in the following, we assume that these are fixed. We use the widely implemented SNAP Bispectrum descriptors [23] with $N_{D}=55$ (see methods), giving a potential energy
@@
where ${{\bm{\Theta}}}\in\mathbb{R}^{N_{D}}$ is the potential parameter vector and ${\bf D}(\tilde{\bf X},{\bf C})\in\mathbb{R}^{N_{D}}$ is the total descriptor vector. The advantage of the linear functional form is that $\nabla_{{{\bm{\Theta}}}}U={\bf D}$ and $\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}U=\nabla_{\tilde{\bf X}}{\bf D}$, required for the solution of equation (5), are readily evaluated without numerical or automatic differentiation schemes.
We use DFT training data for tungsten from Goryaeva et al.[8], generating 100 samples ${{\bm{\Theta}}}_{m}$ from a parameter distribution using the approach described in [9]. The potential samples are suitable for multi-scale uncertainty quantification, for which a more detailed study will be presented elsewhere. Here, we are primarily interested in using the variations of parameter samples away from the reference potential $\bar{{{\bm{\Theta}}}}$ to test our implicit expansion method. To this end, we define an additional ‘perturbation magnitude’ $\lambda$ and generate samples with
@@
such that $\lambda=0$ corresponds to the reference potential $\bar{{{\bm{\Theta}}}}$ and $\lambda=1$ corresponds to the original sample ${{\bm{\Theta}}}_{m}$. We then generated very large (and thus often unphysical) perturbations with $\lambda\in[-25,25]$, with a step $\Delta\lambda=0.2$, truncating only when the bcc lattice became unstable. This yielded a total ensemble of around $20000$ stable potentials. For each potential, we calculated the formation vacancy defect, allowing for relaxation of both structure and volume, meaning that only the full ih+h expansion is expected to be asymptotically exact. The diversity of the resultant dataset allows for a robust test of our implicit Taylor expansions (35) and (7).
Figure 4: Targeting dislocation core reconstruction in tungsten. a Dislocation cluster supercell. The $\sim$2000-atom cell has fixed boundary conditions in the $\langle 110\rangle$ and $\langle 112\rangle$ directions and periodic boundary conditions in the $\langle 111\rangle$ direction with one Burgers vector within the cell, with one ‘substitutional’ atom at the dislocation core, with interaction parameters initially corresponding to tungsten. Around 500 atoms are fixed at the cluster border to stabilize the dislocation core structure. Non-bcc atoms are identified using the OVITO package [31]. b Maximal difference of atomic positions between a current and the target configurations during the minimization procedure. c, d Differential displacement maps [32]. The starting configuration for both simulations is the easy core and the target configurations are the hard core (c) and split core (e). The transient dislocation core positions as functions of the minimization iteration are presented with lines starting with empty and ending at solid triangles as initial and final core configurations. d, f Atomic configurations around the dislocation core with hard code (d) and split core (f) target configurations.
Figures 3a and 3b illustrate this approach. As expected, this form of perturbation applied to all stable potentials produced a wide range of very strong perturbations. Fig. 3c shows the formation volume variation across the samples as a function of perturbation magnitude $\lambda$. Both homogeneous and inhomogeneous expansions provide a nearly-perfect predictions of vacancy formation volume. However, for the formation energy, the inhomogeneous approach is notably more accurate, with errors under 2% across a wide (3 eV) range of formation energies (see Fig. 3d). This indicates that whilst the efficient homogeneous expansion offers a very useful prediction of energy and volume changes, the inhomogeneous term allows for accurate prediction at small to moderate perturbations. This asymptotic accuracy is particularly important when using the implicit derivative to solve inverse problems, which we discuss in the next section.
@@ II.8 Implicit loss minimization applied to solute-induced dislocation core reconstruction
In this final section we employ the implicit derivative concept to solve a challenging inverse problem: with a starting potential ${{\bm{\Theta}}}$, and some stationary configurations ${\bf X}^{*}_{{\bm{\Theta}}}$, we search for the potential parameters that stabilize a structure as close as possible to some target configuration ${\bf X}^{*}$. As discussed in section II.5 this has application for potential fine-tuning, as we demonstrate below. More generally, the ability to find parameters which yield certain desired structures represents a first step towards a range of inverse design strategies, in particular, given the ability of emerging foundation models to smoothly interpolate across chemical space [15].
To demonstrate minimization of the implicit loss function 9, we selected a computationally challenging system of a $\sim$2000-atom tungsten disk with a $\langle 111\rangle/2$ screw dislocation along the disk axis. As detailed in [4], the outer layers of atoms in the disk are fixed to displacement from elasticity theory and we impose periodic boundary conditions along the dislocation line direction. Using initial potential parameters ${{\bm{\Theta}}}_{0}$ for W from Goryaeva et al. [8], the dislocation core relaxes into the ‘easy’ core structure in agreement with ab initio, as illustrated in Fig. 4a. For the inverse problem, we assign one ‘alchemical’ atom (the red atom in Fig. 4a) with it’s own independent set of parameters ${{\bm{\Theta}}}$, initially set to ${{\bm{\Theta}}}_{0}$. We use a simple form for the linear multi-specie potential, detailed in the methods (IV.2). Modifying the alchemical potential parameters ${{\bm{\Theta}}}$ whilst keeping ${{\bm{\Theta}}}_{0}$ fixed, we aim to stabilize the ‘hard’ and ‘split’ core structures which are unstable for W [8]. We generated the target core configurations with the matscipy.dislocation [33] Python package. As the target structures are for pure, single-element W we do not expect the structure induced by the alchemical solute to give an exact match, but we can monitor the effective dislocation core position using the strain-matching approach detailed in [33]. The implicit loss minimization is achieved through the procedure presented in Algorithm II.8. The loss gradient is computed according to equation (10). We employ an adaptive step size $h$ as detailed in SM.
@@
Figure 4b shows the maximal deviation of atomic positions at iteration $k$ for two target structures. The minimization error decreases significantly at first steps and saturates at $\sim$10 iterations for both structures. The error does not reach zero, which we attribute to the target configurations being derived from pure tungsten systems, whereas our minimization involves systems that more closely represent substitutional defects. However, the minimization goals are clearly achieved as seen in Figure 4 panels c-f: the hard core (Fig. 4c,d) and split core (Fig. 4e,f) are located at expected positions denoted by solid triangles in Fig. 4c,e.
As a final application, we show how the implicit loss minimization can be used to fine-tune an initial interatomic potential to match ab initio training data. Here, our target is a solute-induced reconstruction of the $\langle 111\rangle/2$ screw dislocation core caused by an interstitial Be atom, using the same disk geometry as described above and shown in Fig. 5a. The target data was generated using the QM/ML simulation method which embeds a DFT region at core as detailed in methods IV.3. As shown in Fig. 5c, we see that Be induces reconstruction to the ‘hard’ core structure, with the Be interstitial sitting at the center of the dislocation.
Using W-Be ab initio training data from Wood et al. [34], we created an initial set of Be interaction parameters ${{\bm{\Theta}}}$ using the same linear multi-species interatomic potential as above (methods IV.2) with W parameters ${{\bm{\Theta}}}_{0}$ set to those from Goryaeva et al. [8]. The relaxed structure using the initial potential fit is shown in Fig. 5b. It can be seen that in contrast to the QM/ML target, the dislocation core remains in the ‘easy’ configuration, with the Be atom lying outside of the central core region. Using the procedure described above, we performed implicit loss minimization using this initial relaxed configuration. As shown in the SM, the implicit loss achieved near-perfect reproduction of the core reconstruction in around 20 iterations. Future work will investigate more sophisticated minimization schemes than the simple gradient descent used here.
@@ III Discussion
In this paper, we have investigated the use of the implicit derivative of the relaxed atomic structure with interatomic potential parameters, giving a first-order implicit Taylor expansion for the relaxed structures and second order for relaxed energies. We detailed how the implicit derivative could be calculated using dense linear algebra, requiring Hessian evaluation, automatic differentiation, or a Hessian-free linear operator approach, which reduces to a constrained minimization in LAMMPS, allowing application to arbitrarily large systems.
@@
The implicit expansion enables very efficient forward propagation of parameter uncertainties to simulation results, including the effect of geometry relaxation, essential to capture changes in structure such as strain. This was demonstrated on simple classical models and machine learning models for pure W [8]. The implicit expansion was able to capture a wide range of changes in energy and structure, far beyond typical variations associated with potential parameter uncertainty. A forthcoming publication will demonstrate the implicit expansion in a wide-ranging uncertainty quantification study. Beyond uncertainty quantification, our results show that the implicit expansion can also be used to rapidly explore the parameter space of high dimensional interatomic models, permitting parametric studies that would be intractable with standard methods. In future work, we will explore how the implicit expansion can be used in a correlative study of defect structures and implications for both uncertainty quantification and materials design goals.
@@
In addition to the forward propagation enabled by the implicit expansion, we also investigated the use of the implicit derivative in backpropagation of structural changes to parameter changes. Our exploratory applications focused on solute-induced dislocation core reconstruction in W [4], a key feature in understanding plasticity and irradiation damage in bcc metals [35]. We showed how the implicit derivative could be used to ‘fine-tune’ parameters from an initial fit against training data for WBe [34], in order to stabilize the structure seen in ab initio calculations [4].
@@
Figure 5: ‘Fine-tuning’ Be-induced dislocation core reconstruction in tungsten. a Dislocation cluster supercell similar to that shown in figure 4, but with a Be interstitial in the center of the cluster and no substitution. b, c, d differential displacement maps [32] with core and interstitial positions corresponding to the initial SNAP potential (panel b), target configuration obtained with QM-ML (methods IV.3, panel c), and the minimum of the ‘fine-tuned’ potential (panel d).
@@
In a first effort towards targeted ‘alchemical’ design applications, we used the implicit derivative to find substitutional solute interaction parameters which stabilized ‘hard’ or ‘split’ dislocation cores in pure W. The success of this effort extends the scope of alchemical machine learning to large-scale simulations essential to for mechanistic studies of e.g. plasticity. We anticipate that both approaches will gain ever-increasing application with the advent of general-purpose ‘universal’ machine learning potentials, which will be a focus of future efforts.

### v2

@@ II.5 Backpropagation of parameter variations for inverse design problems
Inverse design aims to produce materials with specified (desirable) properties through inverting structure-property relationships. However, this typically requires high-throughput searches, which necessarily cannot afford to perform atomistic simulations of e.g. defect structures and mechanisms to predict mechanical properties. The implicit derivative can be used to make a first step toward gradient-led inverse design, allowing us to find interatomic parameters ${{\bm{\Theta}}}$ that stabilise some atomic configuration ${\bf X}^{*}_{\rm ref}$ observed in an ab initio-accurate simulation e.g. using the density-functional theory (DFT) [4]. We define an implicit loss function:
$L({{\bm{\Theta}}})=\frac{1}{2}\|{\bf X}^{*}_{{{\bm{\Theta}}}}-{\bf X}^{*}_{\rm ref}\|^{2}.$ (9)
To effectively minimize the loss, the loss gradient is required, and the implicit derivative is necessary to compute the derivative of the loss via the chain rule:
$\nabla_{{{\bm{\Theta}}}}L({{\bm{\Theta}}})=\nabla_{{{\bm{\Theta}}}}{\bf X}^{*}_{{{\bm{\Theta}}}}\cdot\big{[}{\bf X}^{*}_{{{\bm{\Theta}}}}-{\bf X}^{*}_{\rm ref}\big{]}\in\mathbb{R}^{N_{D}}.$ (10)
An immediate application of the implicit loss, equation (9), is the ability to ‘fine-tune’, or retrain [4], interatomic potentials to reproduce important DFT minima. Fine-tuning has gained increasing interest following the rise of ‘universal’ machine learning potentials [18]. In practice, one typically includes the loss against the original training database to regularize the fine-tuning fit. However, we have found that minimizing (9) in practice produces only very small perturbations to the final potential.
The ability to fine-tune interactions is of particular relevance for low energy structures such as dislocation lines [8, 4], which typically require careful weighting in potential fitting [18]. In section II.8, we use implicit loss minimization to find solute substitutions which induce ‘hard’ screw dislocation core reconstruction in tungsten, and fine-tune a tungsten-beryllium potential to correctly reproduce ab initio observations.
@@ II.6 Prediction of the total relaxed energy in classical potentials with automatic differentiation
The binary Lennard-Jones potential is a classical model for nanoclusters and glassy systems [5]. The model is defined by six parameters, ${{\bm{\Theta}}}=[\epsilon^{\rm LJ}_{\rm AA},\epsilon^{\rm LJ}_{\rm AB},\epsilon^{\rm LJ}_{\rm BB},\sigma_{\rm AA},\sigma_{\rm AB},\sigma_{\rm BB}]$, with a total energy
@@
To simplify the problem, we set $\epsilon^{\rm LJ}_{\rm AA}=\epsilon^{\rm LJ}_{\rm AB}=\epsilon^{\rm LJ}_{\rm BB}$ and $\sigma_{\rm AA}=\sigma_{\rm BB}=1$, leaving $\Theta=\sigma_{\rm AB}$ as the only varying parameter in this example. When $\sigma_{\mathrm{AB}}=1$, all atoms are identical and the system is a unary fcc lattice; we additionally remove one atom to form a vacancy and promote additional deformation. When $\sigma_{\mathrm{AB}}\neq 1$ the system becomes a random fcc binary alloy, with lattice distortion in the bulk and around the vacancy (see Fig. 2a and 2b). Figure 2c shows the inhomogeneous implicit derivative around $\sigma_{AB}=1$ that gives an excellent prediction of the total energy and lattice distortion for $\sigma_{AB}\in[0.95,1.05]$, with mild disagreement as $|\sigma_{AB}-1|$ grows. At constant volume, both constant and homogeneous approximations are equivalent and predict no structural change, with significantly higher errors. In the remainder of this paper, we focus on modern machine learning potentials.
@@ II.7 Prediction of defect formation energies and volumes with machine learning potentials
Almost all modern interatomic models use high-dimensional regression techniques from the machine learning community [35, 36, 8, 18]. A common first step is to represent atomic environments as $N_{D}$ per-atom descriptor functions $D_{l}(\tilde{\bf X},{\bf C},i)$, $l\in[1,N_{D}]$ which describe the atomic environment around an atom of index $i$. In practice, the descriptor functions also have species-dependent hyperparameters which must also be tuned, but in the following, we assume that these are fixed. We use the widely implemented SNAP Bispectrum descriptors [29] with $N_{D}=55$ (see methods, section IV.3, for multi-specie generalization), giving a potential energy
@@
where ${{\bm{\Theta}}}\in\mathbb{R}^{N_{D}}$ is the potential parameter vector and ${\bf D}(\tilde{\bf X},{\bf C})\in\mathbb{R}^{N_{D}}$ is the total descriptor vector. The advantage of the linear functional form is that $\nabla_{{{\bm{\Theta}}}}U={\bf D}$ and $\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}U=\nabla_{\tilde{\bf X}}{\bf D}$, required for the solution of equation (5), are readily evaluated without any numerical or automatic differentiation schemes.
For multi-scale uncertainty quantification study, we use DFT training data for tungsten from Goryaeva et al. [8] and generate 100 samples ${{\bm{\Theta}}}_{m}$ ($m$ is a sample index) from a parameter distribution using the approach described in [9]. We are primarily interested in using the variations of parameter samples away from the reference potential $\bar{{{\bm{\Theta}}}}$ to test our implicit expansion method. To this end, we define an additional ‘perturbation magnitude’ $\lambda$ and generate samples with
@@
such that $\lambda=0$ corresponds to the reference potential $\bar{{{\bm{\Theta}}}}$ and $\lambda=1$ corresponds to the original sample ${{\bm{\Theta}}}_{m}$. We then generate very large (and thus sometimes unphysical) perturbations with $\lambda\in[-25,25]$, with a step $\Delta\lambda=0.2$, truncating only when the bcc lattice became unstable. This yielded a total ensemble of around $20000$ stable potentials. For each potential, we calculated the formation vacancy defect [37], allowing for relaxation of both structure and volume, meaning that only the full ih+h expansion is expected to be asymptotically exact. The diversity of the resultant dataset allows for a robust test of our implicit Taylor expansions (36) and (7).
Figure 4: Targeting dislocation core reconstruction in tungsten. a Dislocation cluster supercell. The $\sim$2000-atom cell has fixed boundary conditions in the $\langle 110\rangle$ and $\langle 112\rangle$ directions and periodic boundary conditions in the $\langle 111\rangle$ direction with one Burgers vector within the cell. The ‘substitutional’ atom at the dislocation core (in red), initially has interaction parameters corresponding to tungsten, which vary during the loss minimization procedure. Around 500 atoms are fixed at the cluster border to stabilize the dislocation core structure [4]. Non-bcc atoms are identified using the OVITO package [38]. b Maximal difference of atomic positions between a current and the target configurations during the loss minimization procedure for target configurations set to ‘hard’ and ‘split’ cores. c, e Differential displacement maps [39] corresponding to final configurations of the minimization. The starting configuration for both simulations is the ‘easy’ core and the target configurations are the ‘hard’ core (c) and ‘split’ core (e). Transient dislocation core positions during the minimization are presented with lines starting at empty and ending at solid triangles, that depict initial and final core positions. d, f Atomic configurations around the dislocation core with ‘hard’ core (d) and ‘split’ core (f) target configurations.
Figures 3a and 3b illustrate this approach. As expected, this form of perturbation applied to all stable potentials produced a wide range of very strong perturbations. Fig. 3c shows the formation volume error across the samples as a function of perturbation magnitude $\lambda$. Both homogeneous (h) and full (h+ih) expansions provide a nearly-perfect predictions of vacancy formation volume. However, for the formation energy, the full approach is notably more accurate, with errors under 2% across a wide (3 eV) range of formation energies (see Fig. 3d). This indicates that whilst the efficient homogeneous expansion offers a useful prediction of energy and volume changes, the inhomogeneous term allows for most accurate prediction at small to moderate perturbations. This asymptotic accuracy is particularly important when using the implicit derivative to solve inverse problems, which we discuss in the next section.
@@ II.8 Implicit loss minimization applied to solute-induced dislocation core reconstruction
In this final section we employ the implicit derivative concept to solve a challenging inverse problem: with a starting potential ${{\bm{\Theta}}}$, and some stationary configurations ${\bf X}^{*}_{{\bm{\Theta}}}$, we search for the potential parameters that stabilize a structure as close as possible to some target configuration ${\bf X}^{*}_{\rm ref}$. The ability to find parameters that yield certain desired structures represents a first step towards a range of inverse design strategies, in particular, given the ability of emerging universal potential models to smoothly interpolate across chemical space [20].
To demonstrate minimization of the implicit loss function (equation (9)), we selected a computationally challenging system of a $\sim$2000-atom tungsten disk with a $\langle 111\rangle/2$ screw dislocation along the disk axis, as illustrated on Fig. 4a. Using initial potential parameters ${{\bm{\Theta}}}_{0}$ for W from Goryaeva et al. [8], the dislocation core relaxes into the ‘easy’ core structure in agreement with ab initio. For the inverse problem, we assign one ‘alchemical’ atom (the red atom in Fig. 4a) with it’s own independent set of parameters ${{\bm{\Theta}}}$, initially set to ${{\bm{\Theta}}}_{0}$. We use a simple form for the linear multi-specie potential, detailed in the methods, section IV.2. Modifying the alchemical potential parameters ${{\bm{\Theta}}}$ whilst keeping ${{\bm{\Theta}}}_{0}$ fixed, we aim to stabilize the ‘hard’ and ‘split’ core structures, which are unstable for W [40, 8]. We monitor the effective dislocation core position during the minimization using the strain-matching approach detailed in [4]. The implicit loss minimization is achieved through the procedure presented in Algorithm II.8, where the loss gradient is computed according to equation (10). We employ an adaptive step size $h$ as detailed in supplementary material.
@@
Figure 4b shows the maximal deviation of atomic positions at iteration $k$ for two target structures. The minimization error decreases significantly at first steps and saturates at $\sim$10 iterations for both structures. The error does not reach zero, which we attribute to the target configurations being derived from pure tungsten systems, whereas our minimization involves systems that more closely represent substitutional defects. However, the minimization goals are clearly achieved as seen in Figure 4 panels c-f: the ‘hard’ core (Fig. 4c,d) and ‘split’ core (Fig. 4e,f) are located at expected positions denoted by solid triangles.
As a final application, we show how the implicit loss minimization can be used to fine-tune an initial interatomic potential to match ab initio training data. Here, our target is a solute-induced reconstruction of the $\langle 111\rangle/2$ screw dislocation core caused by an interstitial Be atom, using the same disk geometry as described above and shown in Fig. 5a. The target data was generated using the QM/ML simulation method which embeds a DFT region at core as detailed in methods IV.3. As shown in Fig. 5b, we see that Be induces reconstruction to the ‘hard’ core structure, with the Be interstitial sitting at the center of the dislocation.
Using W-Be ab initio training data from Wood et al. [41], we created an initial set of Be interaction parameters ${{\bm{\Theta}}}$ using the same linear multi-species interatomic potential as above (see methods, section IV.2) with W parameters ${{\bm{\Theta}}}_{0}$ set to those from Goryaeva et al. [8]. The relaxed structure using the initial potential fit is shown in Fig. 5c. It can be seen that in contrast to the QM/ML target, the dislocation core remains in the ‘easy’ configuration, with the Be atom lying outside of the central core region. Using the procedure described above, we performed implicit loss minimization using this initial relaxed configuration. The implicit loss achieved near-perfect reproduction of the core reconstruction in around 20 iterations (see Fig. 5d and supplementary material).
@@ III Discussion
In this paper, we have investigated the use of the implicit derivative of the relaxed atomic structure with interatomic potential parameters, giving a first-order implicit Taylor expansion for the relaxed structures and second-order for relaxed energies. We detailed how the implicit derivative could be calculated using dense linear algebra, requiring Hessian evaluation, automatic differentiation, or a Hessian-free linear operator approach, which reduces to a constrained minimization in LAMMPS, allowing application to arbitrarily large systems.
@@
The implicit expansion enables very efficient forward propagation of parameter uncertainties to simulation results, including the effect of geometry relaxation, essential to capture changes in structure. This was demonstrated on simple classical models and machine learning models for pure tungsten. The implicit expansion was able to capture a wide range of changes in energy and structure, far beyond typical variations associated with potential parameter uncertainty. Beyond uncertainty quantification, our results show that the implicit expansion can also be used to rapidly explore the parameter space of high dimensional interatomic models, permitting parametric studies that would be intractable with standard methods. In future work, we will explore how the implicit expansion can be used in a correlative study of defect structures and implications for both uncertainty quantification and materials design goals.
@@
In addition to the forward propagation enabled by the implicit expansion, we also investigated the use of the implicit derivative in backpropagation of structural changes to changes in potential parameters. Our exploratory applications focused on solute-induced dislocation core reconstruction in tungsten, a key feature in understanding plasticity and irradiation damage in bcc metals [42]. We showed how the implicit derivative could be used to ‘fine-tune’ parameters from an initial fit against training data for tungsten-beryllium, in order to stabilize the structure seen in ab initio calculations.
@@
Figure 5: ‘Fine-tuning’ Be-induced dislocation core reconstruction in tungsten. a Dislocation cluster supercell similar to that shown in figure 4, but with a Be interstitial in the center of the cluster and no substitution. b, c, d differential displacement maps [39] with dislocation core and interstitial positions corresponding to the QM-ML configuration (panel b), initial SNAP potential (panel c), and the minimum of the ‘fine-tuned’ potential (panel d).
@@
In a first effort towards targeted ‘alchemical’ design applications, we used the implicit derivative to find substitutional solute interaction parameters which stabilized ‘hard’ or ‘split’ dislocation cores in pure tungsten. The success of this effort extends the scope of alchemical machine learning to large-scale simulations essential for mechanistic studies of e.g. plasticity. We anticipate that both approaches will gain ever-increasing application with the advent of general-purpose ‘universal’ machine learning potentials, which will be a focus of future efforts.